# SemantiCodec: An Ultra Low Bitrate Semantic Audio Codec for General Sound

### v1

@@ Abstract
Large language models (LLMs) have significantly advanced audio processing through audio codecs that convert audio into discrete tokens, enabling the application of language modelling techniques to audio data. However, traditional codecs often operate at high bitrates or within narrow domains such as speech and lack the semantic clues required for efficient language modelling. Addressing these challenges, we introduce SemantiCodec, a novel codec designed to compress audio into fewer than a hundred tokens per second across diverse audio types, including speech, general audio, and music, without compromising quality. SemantiCodec features a dual-encoder architecture: a semantic encoder using a self-supervised AudioMAE, discretized using k-means clustering on extensive audio data, and an acoustic encoder to capture the remaining details. The semantic and acoustic encoder output are used to reconstruct audio via a diffusion-model-based decoder. SemantiCodec is presented in three variants with token rates of $25$, $50$, and $100$ per second, supporting a range of ultra-low bit rates between $0.31$ kbps and $1.43$ kbps. Experimental results demonstrate that SemantiCodec significantly outperforms the state-of-the-art Descript codec on reconstruction quality. Our results also suggest that SemantiCodec contains significantly richer semantic information than all evaluated audio codecs, even at significantly lower bitrates. Our code and demos are available at https://haoheliu.github.io/SemantiCodec/.
@@ I Introduction
Audio codecs are used for encoding and decoding digital audio for efficient telecommunications and broadcasting [1]. Traditional audio codec encoders compress audio by discarding inaudible details to reduce storage and transmission demands [1]. The degree of compression is typically assessed by the bitrate, indicating the amount of data, in bits per second, used to represent the audio signal, with commonly used bitrate ranges such as $128$ kbps to $320$ kbps for MP3 [2] and $6$ kbps to $510$ kbps with the Opus codec [3].
With the introduction of deep learning, audio codecs have significantly evolved with better audio quality and bitrate efficiency [4]. These cutting-edge codecs utilize vector quantization [5] to learn compact codebooks, whose indices are transmitted instead of raw audio data. The sequence of transmitted indices is also referred to as the token sequence. Unlike traditional audio codecs, neural audio codecs typically operate at lower bit rates while maintaining similar audio quality. For instance, Encodec [6] achieves compression at multiple bitrates between $1.5$ kbps and $24$ kbps, the Descript codec [7] operates at $8$ kbps and $16$ kbps, and the HiFi-Codec [8] pushes the boundaries further by reducing the bitrate to $2$ kbps with acceptable quality [8].
@@
Figure 1: Comparison between SemantiCodec and state-of-the-art codecs. Larger values on the horizontal and vertical axis indicate better reconstruction quality and semantic information, respectively. The size of circles indicates bitrates, where smaller ones denote lower bitrates.
@@
Despite the advancement in audio language modelling, the token rate of audio codecs has become a growing concern. For example, the token rate for a $6$ kbps Descript codec is $600$ per second. The auto-regressive (AR) nature of token generation means that the inference time of an audio language model scales with the length of codec tokens, posing challenges not only in computational efficiency but also in model training, where longer sequences demand more computational resources. Additionally, longer sequences may lead to challenges on long-term dependencies. Studies on long-context language models [18] indicate that language models do not effectively leverage long-term context, often superficially utilising it. While low-bitrate audio codecs are available, like the $1.5$ kbps version Encodec [6], their reconstruction quality, with strong artefacts introduced, often falls short of production standards. This situation underscores a crucial trade-off between efficiency and quality in audio codec development, highlighting the need for codecs that can achieve high-quality audio reconstruction at low bitrates.
@@
The presence and richness of semantic information within token sequences play a crucial role in the learning of language models [19, 20]. Support for this assertion also comes from findings by Toraman et al. [10], who demonstrated that on six text processing tasks tokenizers operating at a higher level of granularity, such as byte pair encoding [21], could substantially outperform character-level tokenizers, which often require the model to expand more capacity for understanding. Despite these insights on the importance of semantic information in the tokens, our initial investigations reveal that existing audio codecs, already employed in audio language modelling, fall short of capturing adequate semantic information, even at relatively high bitrate settings. For instance, when employing the latent encodings of the $6.0$ kbps Descript codec [7] for classification tasks across seven benchmarks in the HEAR benchmark [22], the average accuracy is only $33\%$, as demonstrated in Section V-B. In contrast, without fine-tuning, a self-supervised pretrained AudioMAE encoder [23] results in a significantly higher accuracy of $65\%$. The classification accuracy is the indicator of semantic richness within codecs, which is crucial for audio language modelling. Moreover, our analysis indicates that using just the first vector quantization layer of the $6.0$ kbps Descript codec, which is often considered to be the most important layer and selected for audio language modelling [15], the classification accuracy drops even further to an average of $24\%$. This deficiency in capturing and encoding rich semantic information in audio codecs can potentially hinder the performance of audio language models.
@@
In this paper, we introduce SemantiCodec, a novel audio codec designed to tackle the issues of excessive token rate and insufficient semantic encoding in current audio codecs. SemantiCodec exhibits richer semantic information and similar reconstruction quality with lower bitrates than previous codecs, as illustrated in Figure 1. Our approach uses the strong generative capabilities of diffusion models alongside the rich audio representations learned by the self-supervised AudioMAE. SemantiCodec processes mel-spectrograms through two encoders sequentially with two distinct vector quantization (VQ) layers. The first VQ layer is constructed using centroids derived from k-means clustering [24] performed on a large dataset of AudioMAE features, ensuring the capture of semantic information. In contrast, the second layer employs a conventional learnable VQ mechanism, enhancing the audio reconstruction fidelity of SemantiCodec. Our empirical analysis reveals that the semantic content is predominantly encoded within the first VQ layer, contributing over $95\%$ to the classification accuracy. However, integrating the second VQ layer is critical in significantly elevating the quality and intelligibility of audio reconstruction. The concatenated quantized outputs from both VQ layers serve as the conditional input for a latent diffusion model (LDM), which follows the architecture of AudioLDM [25]. The latent diffusion model utilizes the provided conditions to reconstruct high-quality audio after quantization. Our experiment shows SemantiCodec outperforms the state-of-the-art audio codec at a similar bitrate on reconstruction quality and contains significantly better semantic information on audio tokens, leading to improved classification accuracy in audio understanding. In summary, our contributions are listed as follows:
@@
  • We propose SemantiCodec, which leverages strong generative models and the rich feature learnt by self-supervised model for semantic-driven audio encoding and reconstruction.
@@
  • SemantiCodec achieves strong reconstruction performance across general audio types at exceptionally low token rates of $25$, $50$, and $100$ per second, surpassing counterparts operating at significantly higher token rates.
@@
  • Evaluation on audio classification benchmarks demonstrates the significantly richer semantic information in the token of SemantiCodec, even with a single layer of vector quantization, indicating strong potential in future audio language modelling with the SemantiCodec.
@@ II-A Neural Audio Codecs
Traditional audio codecs, as detailed by Valin et al. [26] and Dietz et al. [27], have demonstrated their ability to achieve low latency audio compression across a variety of audio types. However, these traditional approaches often fail to deliver high-quality audio reconstruction at low-bitrate settings, such as $3.0$ kbps. Pioneering deep-learning-based audio codecs, Garbacea et al. [28] introduce the use of vector-quantized variational auto-encoders (VQ-VAEs) for learning neural codecs tailored to speech data. SoundStream [4] proposes a universal codec adaptable to various audio types, incorporating a residual vector quantization (RVQ) module to enhance the quantization process and a generative adversarial network (GAN) [29] to improve the reconstruction quality. Following a similar trajectory, Encodec [6] advances the capabilities of SoundStream by integrating multi-scale discriminators and a loss-balancing strategy for reconstruction training alongside an additional language model to facilitate further compression. HiFi-Codec [8] introduced group-residual vector quantization (GRVQ), a novel approach aimed at reducing the number of codebooks required while preserving the reconstruction quality. Descript codec [7] offers enhancements to Encodec, achieving significantly superior reconstruction performance.
@@ II-C Conditional Audio Generation
The development of generative models [33, 34, 35] has significantly propelled the field of conditional audio generation. Speech generation technologies [36, 37, 38] have evolved to produce speech conditioned on transcriptions and specific styles, such as speaker identity, emotion, and prosody. The generative model has also enabled novel tasks such as binaural sound synthesis [39] and synthetic speech quality refinement [40]. Meanwhile, the generation of music and sound effects has been extended to be conditioned on textual descriptions [12, 25, 41] and visual cues [42, 43], demonstrating the versatility of generative architecture.
Diffusion models [44] stand out for their exceptional generative capabilities in producing diverse and high-quality samples. Diffusion models offer a more tractable training process than GANs, emerging as a preferred choice for many researchers. To address the computational demands associated with training and inference in high-dimensional spaces, latent diffusion models [45] have been introduced. LDMs operate on a lower-dimensional latent space derived from a VAE, significantly reducing computational complexity while maintaining the generative power of traditional diffusion models. This approach has been successfully applied in various conditional audio generation models, including AudioLDM [25], TANGO [46], AudioSR [47] and Make-An-Audio [48], illustrating the flexibility of the diffusion model.
@@ III SemantiCodec
Figure 2: SemantiCodec architecture. For an input audio clip, quantized semantic representation ${\bm{E}}_{s}$ is obtained via a codebook pre-computed using k-means clustering on the AudioMAE embeddings. Then $\mathbf{Y}$ and $\mathbf{{\bm{E}}_{\text{s}}}$ are concatenated and fed to a residual encoder to complement acoustic details, which is discretized to ${\bm{E}}_{a}$ by a vector quantization module. SemantiCodec encoder output ${\bm{E}}$ is obtained by concatenating ${\bm{E}}_{s}$ and ${\bm{E}}_{a}$. A latent diffusion model is trained to generate the original audio clip conditioned on ${\bm{E}}$. The snowflake and fire symbols denote frozen and learnable parameters, respectively.
@@ III-A System Overview
As shown in Figure 2, given an input audio ${\bm{x}}\in\mathbb{R}^{l}$, where $l$ denotes the sample length of audio, we initially transform ${\bm{x}}$ into the mel-spectrogram ${\bm{X}}\in\mathbb{R}^{T\times F}$, with $T$ and $F$ indicating the temporal and frequency dimensions, respectively. Leveraging a pretrained AudioMAE $\mathcal{A}(\cdot)$, we compute the AudioMAE feature $\tilde{{\bm{Y}}}=\mathcal{A}(X)=[\tilde{{\bm{y}}}_{1},\tilde{{\bm{y}}}_{2},\ldots,\tilde{{\bm{y}}}_{L}]\in\mathbb{R}^{L\times E}$, where $L=\frac{TF}{P^{2}}$ denotes the number of patch embedding vectors, and $P$ and $E$ represent the patch size and embedding size of AudioMAE, respectively. Each patch is a distinct, non-overlapping block of the mel-spectrogram processed by the AudioMAE, with multiple patches collectively forming the input to the AudioMAE encoder.
To reduce the number of patch embedding vectors, which directly influence the bitrate after quantization, we aggregate adjacent vectors of $\tilde{{\bm{Y}}}$ into ${\bm{Y}}=[{\bm{y}}_{1},{\bm{y}}_{2},\ldots,{\bm{y}}_{\frac{L}{K}}]\in\mathbb{R}^{\frac{L}{K}\times KE}$, where $K\in\{1,2,4\}$ is the stack factor, yielding ${\bm{y}}_{i}=[\tilde{{\bm{y}}}_{i},\ldots,\tilde{{\bm{y}}}_{i+K-1}]$ for $i\in\{1,2,\ldots,\frac{L}{K}\}$. Following extensive clustering on the vector ${\bm{y}}_{i}$, we derive the semantic codebook $\mathbb{E}_{\text{s}}=[{\bm{e}}_{1},{\bm{e}}_{2},\ldots,{\bm{e}}_{N_{\text{s}}}]$, where $N_{\text{s}}$ denotes the number of entries in the semantic codebook. We refer to this clustering process as semantic clustering.
The stacked feature ${\bm{Y}}$ undergoes initial quantization by $\mathbb{E}_{\text{s}}$ into semantic tokens ${\bm{c}}_{\text{s}}$ and semantic feature ${\bm{E}}_{\text{s}}\in\mathbb{R}^{\frac{L}{K}\times KE}$. Subsequently, we concatenate ${\bm{Y}}$ and ${\bm{E}}_{\text{s}}$, employing an acoustic encoder $\mathcal{F}(\cdot)$ to compute the acoustic feature ${\bm{Y}}_{\text{A}}$, which is then quantized via an acoustic vector quantization layer with entry $\mathbb{E}_{\text{a}}\in\mathbb{R}^{N_{\text{a}}\times E}$, outputting the acoustic tokens ${\bm{c}}_{\text{a}}$ and quantized acoustic feature ${\bm{E}}_{\text{a}}$. The final token for the input audio ${\bm{x}}$ is a merge of semantic and acoustic tokens: ${\bm{c}}=[{\bm{c}}_{\text{s}},{\bm{c}}_{\text{a}}]$.

### v2

@@ Abstract
Large language models (LLMs) have significantly advanced audio processing through audio codecs that convert audio into discrete tokens, enabling the application of language modelling techniques to audio data. However, traditional codecs often operate at high bitrates or within narrow domains such as speech and lack the semantic clues required for efficient language modelling. Addressing these challenges, we introduce SemantiCodec, a novel codec designed to compress audio into fewer than a hundred tokens per second across diverse audio types, including speech, general sound, and music, without compromising quality. SemantiCodec features a dual-encoder architecture: a semantic encoder using a self-supervised pre-trained Audio Masked Autoencoder (AudioMAE), discretized using k-means clustering on extensive audio data, and an acoustic encoder to capture the remaining details. The semantic and acoustic encoder outputs are used to reconstruct audio via a diffusion-model-based decoder. SemantiCodec is presented in three variants with token rates of $25$, $50$, and $100$ per second, supporting a range of ultra-low bit rates between $0.31$ kbps and $1.40$ kbps. Experimental results demonstrate that SemantiCodec significantly outperforms the state-of-the-art Descript codec on reconstruction quality. Our results also suggest that SemantiCodec contains significantly richer semantic information than all evaluated state-of-the-art audio codecs, even at significantly lower bitrates. Our code and demos are available at https://haoheliu.github.io/SemantiCodec/.
@@ I Introduction
Audio codecs are used for encoding and decoding digital audio for efficient telecommunications and broadcasting [1]. Traditional audio encoders compress audio by discarding inaudible details to reduce storage and transmission demands [1]. The degree of compression is typically assessed by the bitrate, indicating the amount of data, in bits per second, used to represent the audio signal, with commonly used bitrate ranges such as $128$ kbps to $320$ kbps for MP3 [2] and $6$ kbps to $510$ kbps for the Opus codec [3].
With the introduction of deep learning, audio codecs have significantly evolved with better audio quality and bitrate efficiency [4]. These cutting-edge codecs utilize vector quantization (VQ) [5] to learn compact codebooks, whose indices are transmitted instead of raw audio data. The sequence of transmitted indices is also referred to as the token sequence. Unlike traditional audio codecs, neural audio codecs typically operate at lower bit rates while maintaining similar audio quality. For instance, Encodec [6] achieves compression at multiple bitrates between $1.5$ kbps and $24$ kbps, the Descript codec [7] operates at $8$ kbps and $16$ kbps, and the HiFi-Codec pushes the boundaries further by reducing the bitrate to $2$ kbps with acceptable quality [8].
@@
Figure 1: Comparison between SemantiCodec and state-of-the-art codecs. Higher values on the horizontal and vertical axis indicate better reconstruction quality and semantic information, respectively. The size of circles indicates bitrates, where smaller ones denote lower bitrates.
@@
Despite the advancement in audio language modelling, the token rate of audio codecs has become a growing concern. For example, the token rate for a $6$ kbps Descript codec is $600$ per second (using six codebooks of size $1024$). The auto-regressive (AR) nature of token generation means that the inference time of an audio language model scales with the rate of codec tokens, posing challenges not only in computational efficiency but also in model training, where longer sequences demand more computational resources. Additionally, longer sequences may lead to challenges on long-term dependencies. Studies on long-context language models [18] indicate that language models do not effectively leverage long-term context, often superficially utilising it. While low-bitrate audio codecs are available, like the $1.5$ kbps version of Encodec [6], their reconstruction quality, with strong artefacts introduced, often falls short of production standards. This situation underscores a crucial trade-off between efficiency and quality in audio codec development, highlighting the need for codecs that can achieve high-quality audio reconstruction at low bitrates.
@@
The presence and richness of semantic information within token sequences play a crucial role in the learning of language models [19, 20]. Support for this assertion also comes from findings by Toraman et al. [10], who demonstrated that on six text processing tasks tokenizers operating at a higher level of granularity, such as byte pair encoding [21], could substantially outperform character-level tokenizers, which often require the model to expand capacity for understanding. Despite these insights on the importance of semantic information in the tokens, our initial investigations reveal that existing audio codecs, already employed in audio language modelling, fall short of capturing adequate semantic information, even at relatively high bitrate settings. For instance, when employing the latent encodings of the $6.0$ kbps Descript codec [7] for classification tasks across seven benchmarks in the HEAR benchmark [22], the average accuracy is only $33\%$, as demonstrated in Section V-B. In contrast, without fine-tuning, a self-supervised pretrained AudioMAE encoder [23] results in a significantly higher accuracy of $61\%$. The classification accuracy is the indicator of semantic richness within codecs, which is crucial for audio language modelling. Moreover, our analysis indicates that using just the first vector quantization layer of the $6.0$ kbps Descript codec, which is often considered to be the most important layer and selected for audio language modelling [15], the classification accuracy drops even further to an average of $25\%$. This deficiency in capturing and encoding rich semantic information in audio codecs can potentially hinder the performance of audio language models.
@@
In this paper, we introduce SemantiCodec, a novel audio codec designed to tackle the issues of excessive token rate and insufficient semantic encoding in current audio codecs. SemantiCodec exhibits richer semantic information and similar reconstruction quality with lower bitrates than previous codecs, as illustrated in Figure 1. Our approach uses the strong generative capabilities of diffusion models alongside the rich audio representations learned by the self-supervised AudioMAE. SemantiCodec processes mel-spectrograms through two encoders sequentially with two distinct vector quantization (VQ) layers. The first VQ layer is constructed using centroids derived from k-means clustering [24] performed on a large dataset of AudioMAE features, ensuring the capture of semantic information. In contrast, the second layer employs a conventional learnable VQ mechanism, enhancing the audio reconstruction fidelity of SemantiCodec. Our empirical analysis reveals that the semantic content is predominantly encoded within the first VQ layer, contributing over $95\%$ to the classification accuracy. However, integrating the second VQ layer is critical in significantly elevating the quality and intelligibility of the reconstructed audio. The concatenated quantized outputs from both VQ layers serve as the conditional input for a latent diffusion model (LDM), which follows the architecture of AudioLDM [25]. The latent diffusion model utilizes the provided conditions to reconstruct high-quality audio after quantization. Our experiment shows that SemantiCodec, at similar bitrates, outperforms the state-of-the-art audio codecs on reconstruction quality and contains significantly richer semantic information in the audio tokens, leading to improved classification accuracy in audio understanding. In summary, our contributions are listed as follows:
@@
  • We propose SemantiCodec, which leverages strong generative models and the rich features learnt by a self-supervised model for semantic-driven audio encoding and reconstruction.
  • To our knowledge, SemantiCodec is the first work to use a diffusion based decoder in a neural audio codec at the time of writing.
@@
  • SemantiCodec achieves strong reconstruction performance across general sound types at exceptionally low token rates of $25$, $50$, and $100$ per second, surpassing counterparts operating at significantly higher token rates.
@@
  • Evaluation on audio classification benchmarks demonstrates the significantly richer semantic information in the tokens of SemantiCodec, even with a single layer of vector quantization, indicating strong potential in future audio language modelling with the SemantiCodec.
@@ II-A Neural Audio Codecs
Traditional audio codecs, as detailed by Valin et al. [26] and Dietz et al. [27], have demonstrated their ability to achieve low latency audio compression across a variety of audio types. However, these traditional approaches often fail to deliver high-quality audio reconstruction at low-bitrate settings, such as $3.0$ kbps. Pioneering deep-learning-based audio codecs, Garbacea et al. [28] introduce the use of vector-quantized variational auto-encoders (VQ-VAEs) for learning neural codecs tailored to speech data. SoundStream [4] proposes a universal codec adaptable to various audio types, incorporating a residual vector quantization (RVQ) module to enhance the quantization process and a generative adversarial network (GAN) [29] to improve the reconstruction quality. Following a similar trajectory, Encodec [6] advances the capabilities of SoundStream by integrating multi-scale discriminators and a loss-balancing strategy for reconstruction alongside an additional language model to facilitate further compression. HiFi-Codec [8] introduced group-residual vector quantization (GRVQ), a novel approach aimed at reducing the number of codebooks required while preserving the reconstruction quality. Descript codec [7] offers enhancements to Encodec, achieving significantly better reconstruction performance.
@@ II-B Semantic Audio Representation Learning
Figure 2: SemantiCodec architecture. For an input audio clip, quantized semantic representation ${\bm{E}}_{s}$ is obtained via a pre-computed codebook using k-means clustering on the AudioMAE embeddings. Then ${\bm{Y}}$ and ${\bm{E}}_{\text{s}}$ are concatenated and fed to a residual encoder to complement acoustic details, which is discretized to ${\bm{E}}_{a}$ by a vector quantization module. SemantiCodec encoder output ${\bm{E}}$ is obtained by concatenating ${\bm{E}}_{s}$ and ${\bm{E}}_{a}$. A latent diffusion model is trained to generate the original audio clip conditioned on ${\bm{E}}$. The snowflake and fire symbols denote frozen and learnable parameters, respectively.
@@ II-C Conditional Audio Generation
The development of generative models [33, 34, 35] has significantly propelled the field of conditional audio generation. Speech generation technologies [36, 37, 38] have evolved to produce speech conditioned on transcriptions and specific styles, such as speaker identity, emotion, and prosody. The generative model has also enabled novel tasks such as binaural sound synthesis [39] and synthetic speech quality refinement [40]. Meanwhile, the generation of music and sound effects has been extended to be conditioned on textual descriptions [12, 25, 41] and visual cues [42, 43], demonstrating the versatility of generative architectures.
Diffusion models [44] stand out for their exceptional generative capabilities in producing diverse and high-quality samples. Diffusion models offer a more tractable training process than GANs, emerging as a preferred choice for many researchers. To address the computational demands associated with training and inference in high-dimensional spaces, latent diffusion models (LDM) [45] have been introduced. LDMs operate on a lower-dimensional latent space derived from a VAE, significantly reducing computational complexity while maintaining the generative power of traditional diffusion models. This approach has been successfully applied in various conditional audio generation models, including AudioLDM [25], TANGO [46], AudioSR [47] and Make-An-Audio [48], illustrating the flexibility of the diffusion model. In this work, we leverage the strong generative modelling ability of LDMs to build a SemantiCodec decoder.
@@ III-A System Overview
As shown in Figure 2, given an input audio ${\bm{x}}\in\mathbb{R}^{l}$, where $l$ denotes the sample length of audio, we initially transform ${\bm{x}}$ into the mel-spectrogram ${\bm{X}}\in\mathbb{R}^{T\times F}$, with $T$ and $F$ indicating the temporal and frequency dimensions, respectively. Leveraging a pretrained AudioMAE $\mathcal{A}(\cdot)$, we compute the AudioMAE features $\tilde{{\bm{Y}}}=\mathcal{A}(X)=[\tilde{{\bm{y}}}_{1},\tilde{{\bm{y}}}_{2},\ldots,\tilde{{\bm{y}}}_{L}]\in\mathbb{R}^{L\times E}$, where $L=\frac{TF}{P^{2}}$ denotes the number of patch embedding vectors, and $P$ and $E$ represent the patch size and embedding size of AudioMAE, respectively. Each patch is a distinct, non-overlapping block of the mel-spectrogram processed by the AudioMAE, with multiple patches collectively forming the input to the AudioMAE encoder.
To reduce the number of patch embedding vectors, which directly influence the bitrate after quantization, we aggregate adjacent vectors of $\tilde{{\bm{Y}}}$ into ${\bm{Y}}=[{\bm{y}}_{1},{\bm{y}}_{2},\ldots,{\bm{y}}_{\frac{L}{K}}]\in\mathbb{R}^{\frac{L}{K}\times KE}$, where $K\in\{1,2,4\}$ is the stack factor, yielding ${\bm{y}}_{i}=[\tilde{{\bm{y}}}_{iK},\ldots,\tilde{{\bm{y}}}_{(i+1)K-1}]$ for $i\in\{0,1,\ldots,\frac{L}{K}-1\}$. Following extensive clustering on the vector ${\bm{y}}_{i}$, we derive the semantic codebook $\mathbb{E}_{\text{s}}=[{\bm{e}}_{1},{\bm{e}}_{2},\ldots,{\bm{e}}_{N_{\text{s}}}]$, where $N_{\text{s}}$ denotes the number of entries in the semantic codebook. We refer to this clustering process as semantic clustering.
The stacked feature ${\bm{Y}}$ undergoes initial quantization by $\mathbb{E}_{\text{s}}$ into semantic tokens ${\bm{c}}_{\text{s}}$ and semantic feature ${\bm{E}}_{\text{s}}\in\mathbb{R}^{\frac{L}{K}\times KE}$. Subsequently, we concatenate ${\bm{Y}}$ and ${\bm{E}}_{\text{s}}$, employing an acoustic encoder $\mathcal{F}(\cdot)$ to compute the acoustic feature ${\bm{Y}}_{\text{A}}$, which is then quantized via an acoustic vector quantization layer with entries $\mathbb{E}_{\text{a}}\in\mathbb{R}^{N_{\text{a}}\times KE}$, outputting the acoustic tokens ${\bm{c}}_{\text{a}}$ and quantized acoustic feature ${\bm{E}}_{\text{a}}$. The final tokens for the input audio ${\bm{x}}$ is a merge of semantic and acoustic tokens: ${\bm{c}}=[{\bm{c}}_{\text{s}},{\bm{c}}_{\text{a}}]$.